OpenStudy (anonymous):

use logarithmic differentiation to find derivative: (1+sin(x))^(1/x)??

myininaya (myininaya):

\[y=(1+\sin(x))^\frac{1}{x}\] Take natural log of both sides \[\ln(y)=\ln(1+\sin(x))^\frac{1}{x}\] \[\ln(y)=\frac{1}{x} \ln(1+\sin(x))\] now we take derivative of both sides we need to recall the product rule and chain rule \[[\ln(y)]'=[\frac{1}{x}]'(1+\sin(x))+\frac{1}{x}[1+\sin(x)]'\] ---- \[[\ln(y)]'=\frac{y'}{y}\] \[[\frac{1}{x}]'=\frac{-1}{x^2}\] \[[1+\sin(x)]'=0+\cos(x)=\cos(x)\] --------------------------

myininaya (myininaya):

opps i left my natural log off

myininaya (myininaya):

\[[\ln(y)]'=[\frac{1}{x}]'\ln(1+\sin(x)+\frac{1}{x}[\ln(1+\sin(x)]'\]

myininaya (myininaya):

\[[\ln(1+\sin(x)]'=\frac{[1+\sin(x)]'}{1+\sin(x)}=\frac{0+\cos(x)}{1+\sin(x)}=\frac{\cos(x)}{1+\sin(x)}\]

myininaya (myininaya):

\[\frac{y'}{y}=\frac{-1}{x^2} \cdot \ln(1+\sin(x))+\frac{1}{x} \cdot \frac{\cos(x)}{1+\sin(x)}\]

myininaya (myininaya):

multiply y on both sides to find y'

OpenStudy (anonymous):

ok i got it you gave me the start i needed